What is the exponent of x^(2/3)=36?
what is the inequality of 2^(x+7) > 1/3^(2)
what is the exponent of 3^(2x-1)=1/243

Answer by Alan3354(69443)

What is the exponent of x^(2/3)=36?
The exponent of x is 2/3.
--------------------------
what is the inequality of 2^(x+7) > 1/3^(2)
That question makes no sense. The statement 2^(x+7) > 1/3^(2) is an inequality.
---------------------------
what is the exponent of 3^(2x-1)=1/243
The exponent of 3 is 2x-1.
--------------------------------
I suspect you want to solve these ??
What is the exponent of x^(2/3)=36?
Try "solve for x"
x^(2/3) = 36
Take the square root of both sides.
x^(1/3) = 6, -6
Cube both sides
x = 216, -216
----------------
2^(x+7) > 1/3^(2)
2^(x+7) > 1/9
(x+7)*log(2) > log(1/9)
x+7 > log(1/9)/log(2)
x > log(1/9)/log(2) - 7
x > -34.34258656
-----------------
3^(2x-1)=1/243
243 = 3^5
3^(2x-1)=3^(-5)
2x-1 = -5
2x = -4
x = -2